OpenStudy (anonymous):

What are the solutions of 2x^2 - 3 = 29? A) ± 16 B) ± 4 C) ± 8 D) ± 13

OpenStudy (ivettef365):

you can start by adding 3 on both sides

OpenStudy (ivettef365):

so you have 2x^2 -3 +3 = 29 + 3 2x^2 = 32

OpenStudy (ivettef365):

Now divide by 2 and you have 2x^2 = 32 -- --- 2 2 x^2 = 16

OpenStudy (ivettef365):

now you can equal the equation to 0 and you have x^2 - 16 = 0 you have the equation as (a^2-b^2) = (a+b)(a-b)

OpenStudy (ivettef365):

so you have (x-4)(x+4) = 0 so x equal to -4 and 4

OpenStudy (anonymous):

thank you

OpenStudy (ivettef365):

your welcome :) Hope you understood explanation